OpenStudy (anonymous):

convert r=35/(9sin(theta)+73cos(theta)) to cartesian

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

I do not understand. You typed a kind of polar function with theta as a variable and r as outcome. Conversion from polar to cartesian coordinates is: x=rcos(theta), y=rsin(theta)

OpenStudy (anonymous):

i need to make the polar equation into a cartesian equation

OpenStudy (anonymous):

the cartesian equation is in the form y=mx+b

OpenStudy (anonymous):

Remember that: \[ x=r\cos\theta \\ y=r\sin\theta \]

OpenStudy (zehanz):

OK, multiply both sides of the equation with 9sin(theta)+73cos(theta). What do you get?

OpenStudy (zehanz):

@abc123! Have you done what I suggested?

OpenStudy (anonymous):

yup i got it thanks

OpenStudy (zehanz):

YW!